OpenStudy (anonymous):

Show that 2 tan 20 + 4 tan 40 + 8 tan 80 = 9(cot 10-tan 10)

OpenStudy (anonymous):

We need to proceed as follows : First we prove that tan x = (cotx - cot2x) cot x - tan x = cos x/ sin x - sin x/ cos x = (cos^2 x - sin ^2 x) /( sinx * cos x) = cot 2x cotx - tanx = cot2x so tan x = cot x - cot 2x